OpenStudy (anonymous):

How would I factor the function f(x) = x^3 + 2x^2 - 25x - 50?

OpenStudy (anonymous):

Factor by grouping.

OpenStudy (cwrw238):

try x +/- the factors of 50 - or yes animal has a good idea

OpenStudy (anonymous):

\[f(x) = x^3 + 2x^2 - 25x - 50=x^2(x+2)-25(x+2)=(x^2-25)(x+2)\]\[=(x+5)(x-5)(x+2)\]

OpenStudy (anonymous):

Remember the patterns, remember the rules, work carefully. Do math every day.

OpenStudy (anonymous):

Thank you!